What mass of iron is need to react with sulfur in order to produce 96 grams of iron (III) sulfide according to the following equation

2Fe + 3S ---->  Fe2S3



1
Expert's answer
2022-02-09T04:42:04-0500

Solution:

Calculate the moles of iron(III) sulfide (Fe2S3):

The molar mass of Fe2S3 is 207.9 g/mol

Hence,

Moles of Fe2S3 = (96 g Fe2S3) × (1 mol Fe2S3 / 207.9 g Fe2S3) = 0.4618 mol Fe2S3


Balanced chemical equation:

2Fe + 3S → Fe2S3

According to stoichiometry:

2 mol of Fe produces 1 mol of Fe2S3

X mol of Fe produces 0.4618 mol of Fe2S3:

Thus:

Moles of Fe = X = (0.4618 mol Fe2S3) × (2 mol Fe / 1 mol Fe2S3) = 0.9236 mol Fe


Calculate the mass of iron (Fe):

The molar mass of Fe is 55.845 g/mol

Hence,

Mass of Fe = (0.9236 mol Fe) × (55.845 g Fe / 1 mol Fe) = 51.58 g Fe


Answer: 51.58 grams of iron (Fe) are needed.